OpenStudy (anonymous):

Int x^4(x5 - 8)^4 dx [0,5] lost on this one, could use some pointers

OpenStudy (anonymous):

use (x5 - 8)=z and then u will get the result

OpenStudy (anonymous):

then find dz which is the dirivitive right?

OpenStudy (anonymous):

dz=5x^4?

OpenStudy (anonymous):

yup dz=5x^4dx

OpenStudy (anonymous):

even change the limit if u want otherwise u can simplify the integration then bring it back in terms ox and then substitute

OpenStudy (anonymous):

ok I think im missing something...\[\int\limits_{?}^{?}x ^{4}(u ^{4})dx\] then \[1/5 x ^{5}(u ^{4})\]

OpenStudy (anonymous):

do I u sub again for the u^4?

OpenStudy (anonymous):

it is 1/5integral of (u^4) du =(u^5)/25 now bring the answer iback in x and the nsubstitute for x

OpenStudy (anonymous):

\[1/5 x^51/5u^5=1/25 x^5u^5?\]

OpenStudy (anonymous):

sorry, I'm having a hard time with this one

OpenStudy (anonymous):

why u r writing x^5 always it seems u hv not cancelled it with x^4

OpenStudy (anonymous):

so it is just 1/25 (u^5)?

OpenStudy (anonymous):

Sorry, lost again

OpenStudy (anonymous):

can you write it down properly with the integrals?

OpenStudy (anonymous):

using formula?.. : /

OpenStudy (anonymous):

\[\int\limits_{?}^{?}x^4(x^5-8)^4\]

OpenStudy (anonymous):

Evaluate the integral

OpenStudy (anonymous):

Tried to u sub for (), but keep getting lost

OpenStudy (anonymous):

I think I got it...

OpenStudy (anonymous):

\[(x^5-8)^4/25\]

OpenStudy (anonymous):

\[(x^5−8)^5/25\]

OpenStudy (anonymous):

i havent seen those for a few years but you can always find the (a-b)^4 and just simplify it until the end...

OpenStudy (anonymous):

just a lot of work... is it supposed to be easy exercise ?

OpenStudy (anonymous):

not really. Usually are a bit challenging

OpenStudy (anonymous):

I have another that I just need a sanity check on...

OpenStudy (anonymous):

then just solve ιt with the time consuming way

OpenStudy (anonymous):

you will need this: (a-b)^4 = a^4 - 4a^3b+6a^2b^2-4ab^3+b^4

OpenStudy (anonymous):

..and just do the calculations

OpenStudy (anonymous):

\[\int\limits_{?}^{?}1/(16-x^2)^2 dx =(16-x^2)^{-1}+c\]

OpenStudy (anonymous):

u=16-x^2 du=-xdx

OpenStudy (anonymous):

\[-\int\limits_{}^{}u^{-2}du\]

OpenStudy (anonymous):

about the second i think you re right

OpenStudy (anonymous):

Hope so. Thanks